Two cars leave Pima Community College traveling in opposite directions. One car travels at 80 mph and the other at 54 mph. In how many hours will they be 938 miles apart?
Found 2 solutions by stanbon, richard1234:
Answer by stanbon(75887)

Two cars leave Pima Community College traveling in opposite directions. One car travels at 80 mph and the other at 54 mph. In how many hours will they be 938 miles apart?
---
The cars are separating at 80+54 = 134 mph
----
time = distance/rate
time = 938/134 = 7 hours
===========================
Cheers,
Stan H.
=============

Answer by richard1234(7193)

Relative to the first car, the second car travels 134 mph. 938/134 = 7 hr.

Btw, you have a Tucson native here (I've actually taken courses at Pima Comm. College).